OpenStudy (swag):

A ladder rests against the side of a building. The base of the ladder is 3 feet from the base of the wall. The top of the ladder rests at a point 4 feet up the wall. How many feet tall is the ladder?

OpenStudy (aravindg):

First draw a figure :)

terenzreignz (terenzreignz):

Diagram helps  And ring up ol' Pythagoras :D

OpenStudy (anonymous):

Does This Help? W^2 + G^2 = L^2 (W=wall=4, G=ground=3, L=Ladder=?) 4^2 + 3^2 = L^2 16 + 9 = L^2 25 = L^2 So the square root of 25 = L or L = 5 forming a standard 3-4-5 right triangle